OpenStudy (anonymous):

If a rancher sold 1/3 of his cattle and then sold another 1/4 of the original amount, what part of his cattle was sold?

OpenStudy (anonymous):

just add those fractions together

OpenStudy (anonymous):

You are trying to find what parts were sold. you sold 1/3 of the cattle and then 1/4 of the original cattle.

OpenStudy (anonymous):

The fully-developed part of his cattle. Alternatively: 1/3+1/4

OpenStudy (anonymous):

Which is what i just said

OpenStudy (anonymous):

correction 1/7

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

7/12

OpenStudy (anonymous):

To add fractions, you need to find a common denominator, then simply add the numerators. So find a number that both 3 and 4 multiply into, then remember to change the numerator in the same way (e.g. if you multiply the denominator by 5, also multiply the numerator by 5). Then simply add all the numerators together over the newly found common denominator, and simplify if necessary. In this case, 3 and 4 both multiply into 12. 3 multiplies in 4 times, hence 1/3 becomes: (4*1)/(4*3) = 4/12. 4 multiplies in 3 times, hence 1/4 becomes: (3*1)/(3*4) = 3/12. Now simply add the numerators: 3+4 = 7, hence: 7/12.